can you please help me solve this problem??

In a particular small town, the population in 1995 was 1200. Ever since then the population has grown by 4% every year. If conditions stay the same, what is the approximate expected population in the year 2000?

Found 2 solutions by ankor@dixie-net.com, stanbon:
Answer by ankor@dixie-net.com(12692)

In a particular small town, the population in 1995 was 1200. Ever since then
the population has grown by 4% every year. If conditions stay the same, what is the approximate expected population in the year 2000?
:
Use the exponential growth equation:
p = 1200(1.04)^5
p = 1200 * 1.21665
p = 1460 folks in 5 yrs which is the yr 2000

Answer by stanbon(48535)

In a particular small town, the population in 1995 was 1200. Ever since then the population has grown by 4% every year. If conditions stay the same, what is the approximate expected population in the year 2000?
------------------------------
Let 1995 correspond to x = 0.
Equation:
Pop(x) = 0.04x + 1200
------------
IN the year 2000 x will equal 5.
Pop(5) = 1200(1+0.04)^5
Pop(5) = 1200*1.2167 = $1459.99
--------------------------------
Cheers,
Stan H.